OpenStudy (anonymous):

Solve for x: l 3x+19 l = 13

OpenStudy (anonymous):

\[|3x+19|=13\] solve two equations \[3x+19=13\] or \[3x+19=-13 \]

OpenStudy (anonymous):

first one is \[3x+19=13\] \[3x=13-19=-6\] \[x=\frac{-6}{2}=-3\]

OpenStudy (anonymous):

second one is \[3x+19=-13\] \[3x=-13-19=-32\] \[x=-\frac{32}{3}\]

OpenStudy (anonymous):

so x = -2 right?

OpenStudy (anonymous):

yeah my mistake! x = -2 is one solution

OpenStudy (anonymous):

\[3x=-6\] \[x=\frac{-6}{3}=-2\]

OpenStudy (anonymous):

second solution is \[3x=-32\] \[x=-\frac{32}{3}\]

OpenStudy (anonymous):

there are two solutions